OpenStudy (anonymous):

x +6y +6z=35 -x -5y -7z=-35 5x +29y +32z=178 -6x -39y +39z=-228

OpenStudy (anonymous):

seems like one too many equations

OpenStudy (anonymous):

x +6y +6z=35=x=-6y+35-6z

OpenStudy (anonymous):

-x -5y -7z=-35=x=-5y+35-7z

OpenStudy (anonymous):

5x +29y +32z=178=x=-29y/5+178/5-32z/5

OpenStudy (anonymous):

-6x -39y +39z=-228=x=13z/2+38-13y/2